# One Homography is All You Need: IMM-based Joint Homography and Multiple Object State Estimation

### v2

@@ Abstract
journal: Information Fusion
@@ 2.2 Tracking by attention
Instead of the pre-defined anchor boxes used in several popular object detectors, Carion et al. [7] decode object positions from a set of learned position embeddings using a transformer decoder. Meinhardt et al. [41] extend this approach to MOT by using previous detections as queries in subsequent frames, resulting in the Trackformer architecture. Whereas Trackformer works on a frame-by-frame basis, Global Tracking Transformers, as proposed by Zhou et al. [42], encode detections in a sliding temporal window of frames. A transformer decoder then associates new detections with these encodings. Having access to an entire measurement sequence, the deep data association module proposed by Pinto et al. [43] eschews the transformer decoder and relies solely on an encoder to associate detections within the given sequence with one another. End-to-end attention-based methods have recently seen success on MOT leaderboards [44, 45, 46]. However, these methods are computationally expensive, especially considering they often train their own detectors in addition to training the attention-based mechanism in an end-to-end approach.
@@ 4.4.2 The Association Algorithm
Figure 5: Graphical representation of the proposed method. Dynamic and static camera motion models are combined with an IMM filter. During association, the normalised Mahlanbobis distances from this filter and the corresponding BIoU scores are weighted proportionally as described in Section 4.4.2. Since the states of the image and ground plane filters are not mixed, the state transitions between them are depicted with dashed lines.
@@ 6.2 Results on the KITTI dataset
At the time of writing, IMM-JHSE is placed twenty-fourth on the KITTI-car leaderboard [^1] https://www.cvlibs.net/datasets/kitti/eval_tracking_detail.php?result=c3e8939f18cba7b23a2e05a92d4df9e994eaaf91. In addition, all methods that are placed above it are 3D MOT methods that make use of regular 3D measurements, except for RAM [88], which uses a self-supervised encoder that optimises a Markov walk along a space-time graph. Table 6 compares IMM-JHSE and some of these methods. Interestingly, while RAM outperforms IMM-JHSE in the car class, IMM-JHSE outperforms RAM in the pedestrian class. While there is a large difference (3.49 HOTA) between IMM-JHSE and the top-ranked method (BiTrack [19]), it should also be noted that IMM-JHSE uses publicly available detections and the corresponding difference in detection accuracy is 3.64%. Furthermore, BiTrack is an offline method.
@@
However, IMM-JHSE shows the poorest results for the pedestrian class in all metrics except AssA in Table 5. As will be seen in Table 7 and Table 8 for the MOT17 and MOT20 datasets, IMM-JHSE seems to perform poorly with tracks which take up relatively little of the image plane area, particularly if a multitude of such tracks are close to one another. In other words, the merits of IMM-JHSE are more evident when tracking relatively large (in terms of bounding box area) objects that move irregularly or with high acceleration or velocity in the image plane rather than in highly cluttered situations. This may be due to the second association stage, which is designed to account for ambiguous detections that may diverge from that expected by the predicted ground plane state. Yet, suppose multiple tracks are close to one another, and all have ambiguous detections. In that case, it is conceivable that they may be confused with one another since they are likely to have high BIoU scores with detections that belong to other tracks.
@@
Another possibility is that the optimised parameters overfit the validation set on the car class, reducing performance on the pedestrian test set. Indeed, the pattern search algorithm seems prone to getting stuck in local minima based on close observation during optimisation iterations.
@@
Despite the relatively poor performance of IMM-JHSE with the pedestrian class, it is placed third on the KITTI-pedestrian leaderboard at the time of writing, where OC-SORT is the only published method that outperforms it.
@@ 6.3 Results on the MOT datasets
IMM-JHSE also significantly outperforms tracking-by-attention methods on the MOT17 dataset. While the performance of IMM-JHSE may be disappointing, it is clear that the high track densities in the MOT datasets also present a challenge to tracking-by-attention methods.
@@
Since MOT20 deals with even more cluttered environments than MOT17, it is unsurprising that IMM-JHSE achieves the worst results for every metric except AssA on the MOT20 test dataset in Table 8.

### v3

@@ Abstract
journal: Expert Systems with Applications
@@ 2.2 Tracking by attention
Instead of the pre-defined anchor boxes used in several popular object detectors, Carion et al. [7] decode object positions from a set of learned position embeddings using a transformer decoder. Meinhardt et al. [41] extend this approach to MOT by using previous detections as queries in subsequent frames, resulting in the Trackformer architecture. Whereas Trackformer works on a frame-by-frame basis, Global Tracking Transformers, as proposed by Zhou et al. [42], encode detections in a sliding temporal window of frames. A transformer decoder then associates new detections with these encodings. Having access to an entire measurement sequence, the deep data association module proposed by Pinto et al. [43] eschews the transformer decoder and relies solely on an encoder to associate detections within the given sequence with one another. End-to-end attention-based methods have recently succeeded on MOT leaderboards [44, 45, 46]. However, these methods are computationally expensive, especially considering they often train their own detectors in addition to training the attention-based mechanism in an end-to-end approach.
@@ 4.4.2 The Association Algorithm
Figure 5: A graphical representation of the proposed method. Dynamic and static camera motion models are combined with an IMM filter. During association, the normalised Mahlanbobis distances from this filter and the corresponding BIoU scores are weighted proportionally as described in Section 4.4.2. Since the states of the image and ground plane filters are not mixed, the state transitions between them are depicted with dashed lines. The transition probabilities are determined as explained in Section 5.
@@ 6.2 Results on the KITTI dataset
At the time of writing, IMM-JHSE is placed twenty-fourth on the KITTI-car leaderboard [^1] https://www.cvlibs.net/datasets/kitti/eval_tracking_detail.php?result=c3e8939f18cba7b23a2e05a92d4df9e994eaaf91. In addition, all methods that are placed above it are 3D MOT methods that make use of regular 3D measurements, except for RAM [88], which uses a self-supervised encoder that optimises a Markov walk along a space-time graph. Table 6 compares IMM-JHSE and some of these methods. Interestingly, while RAM outperforms IMM-JHSE in the car class, IMM-JHSE outperforms RAM in the pedestrian class. While there is a large difference (3.49 HOTA) between IMM-JHSE and the top-ranked method (BiTrack [19]), it should also be noted that IMM-JHSE uses publicly available detections and the corresponding difference in detection accuracy is 3.64%. Furthermore, BiTrack is an offline method. Considering that IMM-JHSE uses only an initial homography estimate and further relies only on 2D detections, it is remarkable that it compares so well to 3D MOT methods (which use regular 3D measurements).
@@
However; in Table 5, which consists of the top three KITTI-pedestrian methods, IMM-JHSE shows the poorest results for the pedestrian class in all metrics except AssA (though it still performs well; it is ranked second on the online KITTI-pedestrian leaderboard — UCMCTrack’s results do not appear there). As will be seen in Table 7 and Table 8 for the MOT17 and MOT20 datasets, IMM-JHSE performs worse with tracks that take up relatively little of the image plane area, particularly if a multitude of such tracks are close to one another. In other words, the merits of IMM-JHSE are more evident when tracking relatively large (in terms of bounding box area) objects that move irregularly or with high acceleration or velocity in the image plane rather than in highly cluttered situations. This may be due to the second association stage, which is designed to account for ambiguous detections that may diverge from that expected by the predicted ground plane state. Yet, suppose multiple tracks are close to one another, and all have ambiguous detections. In that case, it is conceivable that they may be confused with one another since they are likely to have high BIoU scores with detections that belong to other tracks.
It is also possible that pedestrians are not reliably detected by the PermaTrack detector. Figure 8 shows two such examples. Figure 8 also shows a side-effect of the dynamic noise estimation if the initial position or velocity variance is not set high enough: in both subfigures, there are tracks with relatively large uncertainty ellipses. This is because, for the KITTI dataset, the initial velocity variance is very small (see Table 9). Therefore, when a fast-moving track is initialised, IMM-JHSE incorporates the prediction error into the process noise. In these situations, where tracks do not leave the ground plane, explicitly mixing ground plane and bounding box state estimates may be better.
Figure 8: This figure shows two examples where pedestrians are not detected by the PermaTrack detector.
8 The official on the right-hand side is undetected. 8 The couple in front of the car with ID 17 is undetected.
@@
Another possibility is that the optimised parameters overfit the validation set on the car class, reducing performance on the pedestrian test set. Indeed, the pattern search algorithm seems prone to getting stuck in local minima based on close observation during optimisation iterations. This seems especially likely since the KITTI dataset consists primarily of car labels. Figure 9 shows an example of a sequence with few pedestrians.
Figure 9: An example of the large number of car labels relative to pedestrian labels in the KITTI dataset.
@@
Despite its relatively poor performance with the pedestrian class, IMM-JHSE is placed third on the KITTI-pedestrian leaderboard at the time of writing (including UCMCTrack, which does not appear on the online leaderboard), where OC-SORT is the only published method that outperforms it.
@@ 6.3 Results on the MOT datasets
IMM-JHSE also significantly outperforms tracking-by-attention methods on the MOT17 dataset. It is clear that the high track densities in the MOT datasets also present a challenge to tracking-by-attention methods. Figure 10 shows an example of a relatively high track-density frame from the MOT17 dataset. Notice that the ground plane position uncertainty is very high for a particular track in the track cluster on the left. This follows the discussion in Section 6.2 that partial detections can cause ambiguity in IMM-JHSE.
Figure 10: An example of a high track-density frame from the MOT17 dataset.
@@
Since MOT20 deals with even more cluttered environments than MOT17, it is unsurprising that IMM-JHSE achieves the worst results for every metric except AssA on the MOT20 test dataset in Table 8. In contrast to Figure 10, Figure 11 shows an example of a high track-density frame from the MOT20 dataset.
Figure 11: An example of a high track-density frame from the MOT20 dataset.